The magnetic indyction field due to magnetic north pole of strength 10Am and 5cm away from the pole in weber /metre square

Grade:12

1 Answers

Arun
3 years ago
Magnitude of magnetic induction due to a pole of strength m at a distance r from it is
B = (\mu0 / 4\pi)(m / r2) , \mu0 / 4\pi = 10–7 H.
m = 10 Am, r = 5 cm = 5 x 10–2 m
So, magnitude of magnetic induction, B = (10–7)(10 / 25 x 10–4) = 4 x 10–12 Wb / m2.